# The KSBA moduli space of stable log Calabi–Yau surfaces

### 23 Feb 2024 v1

@@ 8.3. Toric degenerations and intrinsic mirrors
The mirror family of the toric degeneration $\mathcal{X}^{gs}_{\widetilde{\mathscr{P}}}\rightarrow\Delta$ can be computed in terms of theta functions defined using broken lines in a scattering diagram $\mathfrak{D}^{gs}$ on $(kP,\widetilde{\mathscr{P}})^{gs}$, constructed by the Kontsevich-Soibelman algorithm from explicit initial walls coming out of the focus-focus singularities – see [38, §6.3] for the construction of the scattering diagram and [44, Appendix A] for a discussion of theta functions. On the other hand, as reviewed in §5.3, the mirror family of $\mathcal{X}_{\widetilde{\mathscr{P}}}\rightarrow\Delta$ can be computed in terms of theta functions defined using broken lines in the canonical scattering diagram $\mathfrak{D}_{\mathcal{X}_{\widetilde{\mathscr{P}}}}$. Arguing as in [40, §3], replacing the use of [45], which applies to a single log Calabi–Yau surface, by [39, §8], which applies to a normal crossing union of log Calabi–Yau surfaces such as $\mathcal{X}_{\widetilde{\mathscr{P}},0}$, one shows that $\mathfrak{D}^{gs}$ can be naturally realized as a deformation of $\mathfrak{D}_{\mathcal{X}_{\widetilde{\mathscr{P}}}}$ obtained by pushing the focus-focus singularities away from the vertices of the triangulation $\widetilde{\mathscr{P}}$. Moreover, as in [40, Lemma 3.28], there is a natural one-to-one correspondence between broken lines and so a natural identification of the corresponding theta functions. The convergence of the algebra of theta functions for $\hat{\mathcal{Y}}^{gs}_{H,a}\rightarrow\mathrm{Spec}\,\mathbb{C}[\![t]\!]$ then follows from the convergence of the algebra of theta functions for $\mathcal{Y}_{\mathcal{X}_{\widetilde{\mathscr{P}}}}\rightarrow S_{\mathcal{X}_{\widetilde{\mathscr{P}}}}$. ∎
@@ 8.4. The general fiber of the mirror family
The integral affine manifold with singularities $(kP)^{gs}$ is naturally an almost toric base as in [90]. Let $(Y^{sym},D^{sym},L^{sym})\rightarrow(kP)^{gs}$ be the corresponding almost toric fibration given by [90, Corollary 4.2]. This fibration is constructed locally as a toric momentum map fibration away from the focus-focus singularities, and using an explicit local model described in [90, §4.2] around the focus-focus singularities. In particular, $Y^{sm}$ is naturally a differentiable orbifold, with orbifold points in one-to-one correspondence with the vertices of $P$, and locally given by the toric surface singularities with momentum polytopes given by the cones locally describing $P$ near its vertices.
@@
By the Construction 2.10 of $(P,\mathscr{P})$ from a toric model of $(Y,D,L)$, $(kP)^{gs}$ is a Symington polytope for $(Y,D,L^{\otimes k})$, and so, it follows from [90, Corollary 4.2] that $(Y,D,L^{\otimes k})$ admits an almost toric fibration over $(kP)^{gs}$ and so is diffeomorphic as an orbifold to $(Y^{sym},D^{sym},L^{sym})$. ∎
@@
The toric mirror degeneration $(\mathcal{Y}^{gs}_{H,a},\mathcal{D}_{H,a}^{gs},\mathcal{L}^{gs}_{H,a})\rightarrow\mathbb{A}^{1}$ admits a gluing description in terms of local charts – see [38, §6.2.5] for details. We claim that it follows from this description that a general fiber $(Y^{gs}_{t},D^{gs}_{t},L^{gs}_{t})$ admits an almost toric fibration over $(kP)^{gs}$, and so is diffeomorphic as an orbifold to $(Y^{sym},D^{sym},L^{sym})$ by [90, Corollary 4.2]. Indeed, away from the finitely many ordinary double points corresponding to focus-focus singularities of $(kP)^{gs}$, the family $(\mathcal{Y}^{gs}_{H,a},\mathcal{D}_{H,a}^{gs},\mathcal{L}^{gs}_{H,a})\rightarrow\mathbb{A}^{1}$ is toric, and the torus fibration over the smooth part of $(kP)^{gs}$ is obtained by locally toric deformation of the torus fibrations on the toric irreducible components of the central fiber induced by the moment maps. Moreover, each of the ordinary double points admits a neighborhood locally described by the equation
@@
Combining Lemma 8.7 and Lemma 8.8, we obtain that $(Y^{gs}_{t},D^{gs}_{t},L^{gs}_{t})$ and $(Y,D,L^{\otimes k})$ are diffeomorphic as orbifolds. Moreover, the orbifold points of both $(Y^{gs}_{t},D^{gs}_{t},L^{gs}_{t})$ and $(Y,D,L^{\otimes k})$ have Zariski neighborhoods isomorphic to the toric surface singularities with momentum polytopes given by the cones locally describing $P$ near its vertices. Indeed, this follows for $(Y,D,L^{\otimes k})$ from the Construction 2.10 of $(P,\mathscr{P})$ from a toric model of $(Y,D,L)$, and for $(Y^{gs}_{t},D^{gs}_{t},L^{gs}_{t})$ from the explicit gluing construction of the mirror to a toric degenerations by gluing local models – see [38, §6.2.5].
@@
By Theorem 8.6, there exists a fiber of the double mirror $(\mathcal{Y}_{\mathcal{X}_{\widetilde{\mathscr{P}}}},\mathcal{D}_{\mathcal{X}_{\widetilde{\mathscr{P}}}},\mathcal{L}_{\mathcal{X}_{\widetilde{\mathscr{P}}}})\rightarrow S_{\mathcal{X}_{\widetilde{\mathscr{P}}}}$ which is a general fiber of one of the mirror families $(\mathcal{Y}^{gs}_{H,a},\mathcal{D}_{H,a}^{gs},\mathcal{L}^{gs}_{H,a})\rightarrow\mathbb{A}^{1}$. Hence, it follows from Theorem 8.9 that the general fiber of the double mirror $(\mathcal{Y}_{\mathcal{X}_{\widetilde{\mathscr{P}}}},\mathcal{D}_{\mathcal{X}_{\widetilde{\mathscr{P}}}},\mathcal{L}_{\mathcal{X}_{\widetilde{\mathscr{P}}}})\rightarrow S_{\mathcal{X}_{\widetilde{\mathscr{P}}}}$ is deformation equivalent to $(Y,D,L^{\otimes k})$. On the other hand, denoting by $(Y_{t},D_{t},L_{t})$ a general fiber of the double mirror $(\mathcal{Y}_{\mathcal{X}_{\mathscr{P}}},\mathcal{D}_{\mathcal{X}_{\mathscr{P}}},\mathcal{L}_{\mathcal{X}_{\mathscr{P}}})\rightarrow S_{\mathcal{X}_{\widetilde{\mathscr{P}}}}$, if follows from Theorem 8.2 that $(Y_{t},D_{t},L_{t}^{\otimes k})$ is a fiber of the double mirror of $(\mathcal{Y}_{\mathcal{X}_{\widetilde{\mathscr{P}}}},\mathcal{D}_{\mathcal{X}_{\widetilde{\mathscr{P}}}},\mathcal{L}_{\mathcal{X}_{\widetilde{\mathscr{P}}}})\rightarrow S_{\mathcal{X}_{\widetilde{\mathscr{P}}}}$, and so is also deformation equivalent to $(Y,D,L^{\otimes k})$. Finally, as $\mathrm{Pic}(Y)$ is torsion-free by Lemma 2.9, $L_{t}^{\otimes k}\simeq L^{\otimes k}$ implies $L_{t}\simeq L$, and this concludes the proof of Theorem 8.10. ∎
@@ 8.5. The Hacking–Keel–Yu conjecture
Let $(Y,D,L)$ be a polarized log Calabi–Yau surface with a maximal degeneration defined by $(P,\mathscr{P})$ as in §2.4, and let ${\mathcal{X}_{\mathscr{P}}}$ be its semistable mirror. Then, the KSBA moduli space $\mathcal{M}_{(Y,D,L)}$ admits a toric finite cover
$S_{\mathcal{X}_{\mathscr{P}}}^{\mathrm{sec}}\longrightarrow\mathcal{M}_{(Y,D,L)}\,,$
where $S_{\mathcal{X}_{\mathscr{P}}}^{\mathrm{sec}}$ is the toric variety associated to the secondary fan $\mathrm{Sec}(\mathcal{X}_{\mathscr{P}}/\overline{\mathcal{X}_{\mathscr{P}}})$ as in Definition 6.26.
@@
By Theorem 7.13, the toric variety $S_{\mathcal{X}_{\mathscr{P}}}^{\mathrm{sec}}$ is the base of a family of KSBA stable log Calabi–Yau surfaces which by Theorem 8.10 are generically deformation equivalent to $(Y,D,L)$. Hence, by the universal property of $\mathcal{M}_{(Y,D,L)}$ we obtain an induced map $S_{\mathcal{X}_{\mathscr{P}}}^{\mathrm{sec}}\to\mathcal{M}_{(Y,D,L)}$.
@@
To prove that this map is finite, it is enough by [49, Lemma 8.7] to show that no 1-dimensional toric stratum of $S_{\mathcal{X}_{\mathscr{P}}}^{\mathrm{sec}}$ is contracted to a point, that is, that $(\mathcal{Y},\mathcal{D}+\epsilon\,\mathcal{C})$ is a non-constant family of KSBA stable pairs in restriction to every 1-dimensional toric stratum of $S_{\mathcal{X}_{\mathscr{P}}}^{\mathrm{sec}}$. Let $\tau$ be a codimension one cone of the secondary fan $\mathrm{Sec}(\mathcal{X}_{\mathscr{P}}/\overline{\mathcal{X}}_{\mathscr{P}})$, corresponding to a 1-dimensional toric stratum $S_{\tau}$ of $S_{\mathcal{X}_{\mathscr{P}}}^{\mathrm{sec}}$. As the fan $\mathrm{Sec}(\mathcal{X}_{\mathscr{P}}/\overline{\mathcal{X}}_{\mathscr{P}})$ is complete, it contains a maximal cone $\sigma$ containing $\tau$, corresponding to a 0-dimensional toric stratum $S_{\sigma}$ contained in $S_{\tau}$. We will show that the fiber $(Y_{\sigma},D_{\sigma})$ of $(\mathcal{Y},\mathcal{D})$ over $S_{\sigma}$ is not isomorphic to the general fiber $(Y_{\tau},D_{\tau})$ of $(\mathcal{Y},\mathcal{D})$ over $S_{\tau}$.
Consider a maximal cone $\mathrm{Nef}(\mathcal{Z}_{\sigma}/\overline{\mathcal{X}}_{\mathscr{P}})+\sum_{v\in I}\mathbb{R}_{\geq 0}(X^{\prime})^{v}$ of the Mori fan contained in $\sigma$, and with a codimension one face $\mathrm{Nef}(\mathcal{Z}_{\tau}/\overline{\mathcal{X}}_{\mathscr{P}})+\sum_{v\in J}\mathbb{R}_{\geq 0}(X^{\prime})^{v}$ contained in $\tau$. Here, we have a projective crepant resolution $\mathcal{X}^{\prime}\rightarrow\overline{\mathcal{X}}_{\mathscr{P}}$, with central fiber $\mathcal{X}^{\prime}_{0}=\bigcup_{v\in P_{\mathbb{Z}}}(X^{\prime})^{v}$, a birational contraction $\mathcal{X}^{\prime}\rightarrow\mathcal{Z}_{\sigma}$ contracting the divisors $(X^{\prime})^{v}$ with $v\in I$, corresponding to a face $H_{\sigma}$ of $NE(\mathcal{X}^{\prime}/\overline{\mathcal{X}}_{\mathscr{P}})$, and a birational contraction $\mathcal{Z}_{\sigma}\rightarrow\mathcal{Z}_{\tau}$ of relative Picard number one, such that the composition $\mathcal{X}^{\prime}\rightarrow\mathcal{Z}_{\tau}$, corresponding to a face $H_{\tau}$ of $NE(\mathcal{X}^{\prime}/\overline{\mathcal{X}}_{\mathscr{P}})$, contracts the divisor $(X^{\prime})^{v}$ with $v\in J$. By Theorem 7.12, the intersection complex of $(Y_{\sigma},D_{\sigma})$ (resp . $(Y_{\tau},D_{\tau})$) is given by the polyhedral decomposition $\mathscr{P}_{H_{\sigma}}$ (resp . $\mathscr{P}_{H_{\tau}}$) associated to $H_{\sigma}$ (resp . $H_{\tau}$) as in §7.3. In general, we clearly have $I\subset J$. If $I\neq J$, then, as $\mathcal{Z}_{\sigma}\rightarrow\mathcal{Z}_{\tau}$ has relative Picard rank one, $J\setminus I$ consists of a single vertex $v$ of $\mathscr{P}_{H_{\sigma}}$. If the corresponding surface $X^{\prime}_{\sigma}$ in $\mathcal{Z}_{\sigma}$ as a boundary component contracted by $\mathcal{Z}_{\sigma}\rightarrow\mathcal{Z}_{\tau}$, then, an edge of $\mathscr{P}_{H_{\sigma}}$ disappears in $\mathscr{P}_{H_{\tau}}$, and so $Y_{\sigma}$, and $Y_{\tau}$ are different. If no boundary component of $X^{\prime}_{\sigma}$ is contracted, then $X^{\prime}_{\sigma}$ is a $\mathbb{P}^{1}$-bundle over $\mathbb{P}^{1}$, contracted to the base $\mathbb{P}^{1}$, and $\partial X^{\prime}_{\sigma}$ is irreducible, mapping 2:1 over the base. The corresponding local geometry and its effect on the mirror are studied in detail in [49, §8.2], and the upshot is that $Y_{\sigma}$ and $Y_{\tau}$ are still different.

### 30 Mar 2024 v2

@@ 8.3. Toric degenerations and intrinsic mirrors
The mirror family of the toric degeneration $\mathcal{X}^{gs}_{\widetilde{\mathscr{P}}}\rightarrow\Delta$ can be computed in terms of theta functions defined using broken lines in a scattering diagram $\mathfrak{D}^{gs}$ on $(kP,\widetilde{\mathscr{P}})^{gs}$, constructed by the Kontsevich-Soibelman algorithm from explicit initial walls coming out of the focus-focus singularities – see [39, §6.3] for the construction of the scattering diagram and [45, Appendix A] for a discussion of theta functions. On the other hand, as reviewed in §5.3, the mirror family of $\mathcal{X}_{\widetilde{\mathscr{P}}}\rightarrow\Delta$ can be computed in terms of theta functions defined using broken lines in the canonical scattering diagram $\mathfrak{D}_{\mathcal{X}_{\widetilde{\mathscr{P}}}}$. Arguing as in [41, §3], replacing the use of [46], which applies to a single log Calabi–Yau surface, by [40, §8], which applies to a normal crossing union of log Calabi–Yau surfaces such as $\mathcal{X}_{\widetilde{\mathscr{P}},0}$, one shows that $\mathfrak{D}^{gs}$ can be naturally realized as a deformation of $\mathfrak{D}_{\mathcal{X}_{\widetilde{\mathscr{P}}}}$ obtained by pushing the focus-focus singularities away from the vertices of the triangulation $\widetilde{\mathscr{P}}$. Moreover, as in [41, Lemma 3.28], there is a natural one-to-one correspondence between broken lines and so a natural identification of the corresponding theta functions. The convergence of the algebra of theta functions for $\hat{\mathcal{Y}}^{gs}_{H,a}\rightarrow\mathrm{Spec}\,\mathbb{C}[\![t]\!]$ then follows from the convergence of the algebra of theta functions for $\mathcal{Y}_{\mathcal{X}_{\widetilde{\mathscr{P}}}}\rightarrow S_{\mathcal{X}_{\widetilde{\mathscr{P}}}}$. ∎
@@ 8.4. The general fiber of the mirror family
The integral affine manifold with singularities $(kP)^{gs}$ is naturally an almost toric base as in [91]. Let $(Y^{sym},D^{sym},L^{sym})\rightarrow(kP)^{gs}$ be the corresponding almost toric fibration given by [91, Corollary 4.2]. This fibration is constructed locally as a toric momentum map fibration away from the focus-focus singularities, and using an explicit local model described in [91, §4.2] around the focus-focus singularities. In particular, $Y^{sym}$ is naturally a differentiable orbifold, with orbifold points in one-to-one correspondence with the vertices of $P$, and locally given by the toric surface singularities with momentum polytopes given by the cones locally describing $P$ near its vertices.
@@
By the Construction 2.10 of $(P,\mathscr{P})$ from a toric model of $(Y,D,L)$, $(kP)^{gs}$ is a Symington polytope for $(Y,D,L^{\otimes k})$, and so, it follows from [91, Corollary 4.2] that $(Y,D,L^{\otimes k})$ admits an almost toric fibration over $(kP)^{gs}$ and so is diffeomorphic as an orbifold to $(Y^{sym},D^{sym},L^{sym})$. ∎
@@
The toric mirror degeneration $(\mathcal{Y}^{gs}_{H,a},\mathcal{D}_{H,a}^{gs},\mathcal{L}^{gs}_{H,a})\rightarrow\mathbb{A}^{1}$ admits a gluing description in terms of local charts – see [39, §6.2.5] for details. We claim that it follows from this description that a general fiber $(Y^{gs}_{t},D^{gs}_{t},L^{gs}_{t})$ admits an almost toric fibration over $(kP)^{gs}$, and so is diffeomorphic as an orbifold to $(Y^{sym},D^{sym},L^{sym})$ by [91, Corollary 4.2]. Indeed, away from the finitely many ordinary double points corresponding to focus-focus singularities of $(kP)^{gs}$, the family $(\mathcal{Y}^{gs}_{H,a},\mathcal{D}_{H,a}^{gs},\mathcal{L}^{gs}_{H,a})\rightarrow\mathbb{A}^{1}$ is toric, and the torus fibration over the smooth part of $(kP)^{gs}$ is obtained by locally toric deformation of the torus fibrations on the toric irreducible components of the central fiber induced by the moment maps. Moreover, each of the ordinary double points admits a neighborhood locally described by the equation
@@
Combining Lemma 8.7 and Lemma 8.8, we obtain that $(Y^{gs}_{t},D^{gs}_{t},L^{gs}_{t})$ and $(Y,D,L^{\otimes k})$ are diffeomorphic as orbifolds. Moreover, the orbifold points of both $(Y^{gs}_{t},D^{gs}_{t},L^{gs}_{t})$ and $(Y,D,L^{\otimes k})$ have Zariski neighborhoods isomorphic to the toric surface singularities with momentum polytopes given by the cones locally describing $P$ near its vertices. Indeed, this follows for $(Y,D,L^{\otimes k})$ from the Construction 2.10 of $(P,\mathscr{P})$ from a toric model of $(Y,D,L)$, and for $(Y^{gs}_{t},D^{gs}_{t},L^{gs}_{t})$ from the explicit gluing construction of the mirror to a toric degenerations by gluing local models – see [39, §6.2.5].
@@
By Theorem 8.6, there exists a fiber of the double mirror $(\mathcal{Y}_{\mathcal{X}_{\widetilde{\mathscr{P}}}},\mathcal{D}_{\mathcal{X}_{\widetilde{\mathscr{P}}}},\mathcal{L}_{\mathcal{X}_{\widetilde{\mathscr{P}}}})\rightarrow S_{\mathcal{X}_{\widetilde{\mathscr{P}}}}$ which is a general fiber of one of the mirror families $(\mathcal{Y}^{gs}_{H,a},\mathcal{D}_{H,a}^{gs},\mathcal{L}^{gs}_{H,a})\rightarrow\mathbb{A}^{1}$. Hence, it follows from Theorem 8.9 that the general fiber of the double mirror $(\mathcal{Y}_{\mathcal{X}_{\widetilde{\mathscr{P}}}},\mathcal{D}_{\mathcal{X}_{\widetilde{\mathscr{P}}}},\mathcal{L}_{\mathcal{X}_{\widetilde{\mathscr{P}}}})\rightarrow S_{\mathcal{X}_{\widetilde{\mathscr{P}}}}$ is deformation equivalent to $(Y,D,L^{\otimes k})$. On the other hand, denoting by $(Y_{t},D_{t},L_{t})$ a general fiber of the double mirror $(\mathcal{Y}_{\mathcal{X}_{\mathscr{P}}},\mathcal{D}_{\mathcal{X}_{\mathscr{P}}},\mathcal{L}_{\mathcal{X}_{\mathscr{P}}})\rightarrow S_{\mathcal{X}_{\widetilde{\mathscr{P}}}}$, it follows from Theorem 8.2 that $(Y_{t},D_{t},L_{t}^{\otimes k})$ is a fiber of the double mirror of $(\mathcal{Y}_{\mathcal{X}_{\widetilde{\mathscr{P}}}},\mathcal{D}_{\mathcal{X}_{\widetilde{\mathscr{P}}}},\mathcal{L}_{\mathcal{X}_{\widetilde{\mathscr{P}}}})\rightarrow S_{\mathcal{X}_{\widetilde{\mathscr{P}}}}$, and so is also deformation equivalent to $(Y,D,L^{\otimes k})$. Finally, as $\mathrm{Pic}(Y)$ is torsion-free by Lemma 2.9, $L_{t}^{\otimes k}\simeq L^{\otimes k}$ implies $L_{t}\simeq L$, and this concludes the proof of Theorem 8.10. ∎
@@ 8.5. The Hacking–Keel–Yu conjecture
Let $(Y,D,L)$ be a polarized log Calabi–Yau surface with a maximal degeneration defined by $(P,\mathscr{P})$ as in §2.4, and let ${\mathcal{X}_{\mathscr{P}}}$ be its semistable mirror. Then, the coarse KSBA moduli space $M_{(Y,D,L)}$ admits a toric finite cover $S_{\mathcal{X}_{\mathscr{P}}}^{\mathrm{sec}}\longrightarrow\mathcal{M}_{(Y,D,L)}$, where $S_{\mathcal{X}_{\mathscr{P}}}^{\mathrm{sec}}$ is the toric variety associated to the secondary fan $\mathrm{Sec}(\mathcal{X}_{\mathscr{P}}/\overline{\mathcal{X}}_{\mathscr{P}})$ as in Definition 6.26.
@@
By Theorem 7.13, the toric stack $\mathcal{S}^{\mathrm{sec}}_{\mathcal{X}}$ is the base of a family of KSBA stable log Calabi–Yau surfaces which by Theorem 8.10 are generically deformation equivalent to $(Y,D,L)$. Hence, by the universal property of the KSBA moduli space $\mathcal{M}_{(Y,D,L)}$ we obtain a map of stacks $\mathcal{S}^{\mathrm{sec}}_{\mathcal{X}_{\mathscr{P}}}\to\mathcal{M}_{(Y,D,L)}$, and so an induced map $S^{\mathrm{sec}}_{\mathcal{X}_{\mathscr{P}}}\to M_{(Y,D,L)}$ at the level of coarse moduli spaces.
@@
To prove that this map is finite, it is enough by [50, Lemma 8.7] to show that no 1-dimensional toric stratum of $\mathcal{S}^{\mathrm{sec}}_{\mathcal{X}_{\mathscr{P}}}$ is contracted to a point, that is, that $(\mathcal{Y},\mathcal{D}+\epsilon\,\mathcal{C})$ is a non-constant family of KSBA stable pairs in restriction to every 1-dimensional toric stratum of $\mathcal{S}^{\mathrm{sec}}_{\mathcal{X}_{\mathscr{P}}}$. Let $\rho_{0}$ be a codimension one cone of the secondary fan $\mathrm{Sec}(\mathcal{X}_{\mathscr{P}}/\overline{\mathcal{X}}_{\mathscr{P}})$, corresponding to a 1-dimensional toric stratum $S_{\rho_{0}}$ of $\mathcal{S}^{\mathrm{sec}}_{\mathcal{X}_{\mathscr{P}}}$. As the fan $\mathrm{Sec}(\mathcal{X}_{\mathscr{P}}/\overline{\mathcal{X}}_{\mathscr{P}})$ is complete, there are two maximal cones $\rho_{1}$ and $\rho_{2}$ containing $\rho_{0}$, corresponding to the two 0-dimensional toric strata $S_{\rho_{1}}$ and $S_{\rho_{2}}$ contained in $S_{\rho_{0}}$. We will show that at least one of the fibers $(Y_{\rho_{1}},D_{\rho_{1}})$ or $(Y_{\rho_{2}},D_{\rho_{2}})$ of $(\mathcal{Y},\mathcal{D})$ over the points $S_{\rho_{1}}$ and $S_{\rho_{2}}$ is not isomorphic to the general fiber $(Y_{\rho_{0}},D_{\rho_{0}})$ of $(\mathcal{Y},\mathcal{D})$ over $S_{\rho_{0}}$.
Let $\tilde{\rho}_{0}$ be a codimension one cone of the Mori fan containing $\rho_{0}$, and $\tilde{\rho}_{1}$, $\tilde{\rho}_{2}$ be the two maximal dimensional cones of the Mori fan containing $\tilde{\rho}_{0}$ and contained in $\rho_{1}$ and $\rho_{2}$ respectively. For $i=0,1,2$, we have $\tilde{\rho}_{i}=\mathrm{Nef}(\mathcal{Z}_{i}/\overline{\mathcal{X}}_{\mathscr{P}})+\sum_{v\in J_{i}}\mathbb{R}_{\geq 0}[X^{v}]$, where $\mathcal{Z}_{i}\rightarrow\overline{\mathcal{X}}_{\mathscr{P}}$ is a partial crepant resolution, and $X^{v}$ with $v\in J_{i}$ are divisors which are contracted by a crepant resolution of $\mathcal{Z}_{i}$. Moreover, since $\tilde{\rho}_{0}$ is a codimension one face of $\tilde{\rho}_{i}$, we have a birational contraction $\mathcal{Z}_{i}\rightarrow\mathcal{Z}_{0}$, and either this map is of relative Picard rank one and $J_{i}=J_{0}$, or this map is an isomorphism and $J_{i}=J_{0}\sqcup\{V\}$ for some $V\in J_{i}\setminus J_{0}$. The later case cannot happen for both $i=1$ and $i=2$: indeed, if $J_{1}=J_{0}\sqcup\{V\}$, then $[X^{V}]\in\tilde{\rho}_{1}$ and $[X^{V}]\notin\tilde{\rho}_{0}$, and so $[X^{V}]\notin\tilde{\rho}_{2}$.
Hence, without loss of generality, we can assume that $\mathcal{Z}_{1}\rightarrow\mathcal{Z}_{0}$ is of relative Picard rank one and $J_{1}=J_{0}$. Let $\mathcal{X}^{\prime}\rightarrow\mathcal{Z}_{1}$ be a projective crepant resolution. Denote by $H_{1}$ the face of $NE(\mathcal{X}^{\prime}/\overline{\mathcal{X}}_{\mathscr{P}})$ corresponding to $\mathcal{X}^{\prime}\rightarrow\mathcal{Z}_{1}$, and by $H_{0}$ the codimension one face of $H_{1}$ corresponding to $\mathcal{Z}_{1}\rightarrow\mathcal{Z}_{0}$. By Theorem 7.12, the intersection complex of $(Y_{\rho_{1}},D_{\rho_{1}})$ (resp. $(Y_{\rho_{0}},D_{\rho_{0}})$) is given by the polyhedral decomposition $\mathscr{P}_{H_{1}}$ (resp. $\mathscr{P}_{H_{0}}$) associated to $H_{1}$ (resp . $H_{0}$) as in §7.3.